OpenStudy (anonymous):

I just need a check, 5x^(2)-7x+5=0, I see X=no real solutions? Is this right?

OpenStudy (anonymous):

yup, (-7)^2-4(5)(5)=49-100<0

jimthompson5910 (jim_thompson5910):

yep you got it

OpenStudy (anonymous):

\[(7\pm \sqrt{-51})/10\]

OpenStudy (anonymous):

thanks, we aren't talking about i yet.